# COBRA: A Continual Learning Approach to Vision-Brain Understanding

### v1

@@ Abstract
Vision-Brain Understanding (VBU) aims to extract visual information perceived by humans from brain activity recorded through functional Magnetic Resonance Imaging (fMRI). Despite notable advancements in recent years, existing studies in VBU continue to face the challenge of catastrophic forgetting, where models lose knowledge from prior subjects as they adapt to new ones. Addressing continual learning in this field is, therefore, essential. This paper introduces a novel framework called Continual Learning for Vision-Brain (COBRA) to address continual learning in VBU. Our approach includes three novel modules: a Subject Commonality (SC) module, a Prompt-based Subject Specific (PSS) module, and a transformer-based module for fMRI, denoted as MRIFormer module. The SC module captures shared vision-brain patterns across subjects, preserving this knowledge as the model encounters new subjects, thereby reducing the impact of catastrophic forgetting. On the other hand, the PSS module learns unique vision-brain patterns specific to each subject. Finally, the MRIFormer module contains a transformer encoder and decoder that learns the fMRI features for VBU from common and specific patterns. In a continual learning setup, COBRA is trained in new PSS and MRIFormer modules for new subjects, leaving the modules of previous subjects unaffected. As a result, COBRA effectively addresses catastrophic forgetting and achieves state-of-the-art performance in both continual learning and vision-brain reconstruction tasks, surpassing previous methods.
1 Introduction
@@
Recently, vision-brain understanding (VBU) has made substantial progress due to advances in generative models, including GANs [53, 65] and Diffusion models [68, 62, 72, 55, 27, 63]. These approaches allow the reconstruction of increasingly realistic and accurate images from functional Magnetic Resonance Imaging (fMRI) signals. The approaches designed per-model-per-subject, as seen in [68, 62], are required to train a new model for a new subject. It can learn subject-specific but ignore subject-commonality features. To address this limitation, the studies [72, 55, 27, 63] proposed unified models trained on multiple subjects at a time. While these approaches are trained on a close set of subjects, they may still face a degradation in performance on new subjects. Recent efforts [72, 63] have aimed to adapt pre-trained models to new subject data, yet they continue to suffer from performance drops on prior subjects due to catastrophic forgetting [70, 19, 59], where models lose learned knowledge when adapting to new subject shifts. The ideal solution to this problem is to update the entire model with prior and new data. However, given the sensitive nature of brain data and the importance of privacy, retaining previous training data is often infeasible. Moreover, as the number of subjects grows, computational training costs become a concern. Thus, in practice, VBU models need to continually learn from new subjects without requiring retraining on prior data. This paradigm is known as Continual Vision-Brain Understanding (CVBU). Given the limitation in previous brain understanding approaches, we raise critical questions for CVBU: (1) What types of knowledge can the model retain, and (2) what types of knowledge must it learn specifically from new subject data?
@@
Since all subjects share similar patterns of brain activations [13, 9, 15, 74, 16, 17, 14, 23, 22, 20], if the model can understand these patterns, there is no need to learn from the data of new subjects. Consequently, the answer to the first question is to retain the knowledge of common patterns of brain activations learned from prior subjects and apply them to new ones. To achieve this, we propose a novel module designed to capture these patterns. Unlike prior methods [13, 9, 15, 74, 16, 17, 14, 23, 22, 20], our proposed approach can automatically identify a broader set of common patterns.
@@
Contributions of this Work: In summary, this work presents a novel continual learning approach to VBU. The contributions of this work can be summarized as follows. First, we introduce a new Continual Learning for Vision-Brain Understanding (COBRA) framework to the visual-brain understanding problem. To the best of our knowledge, this is the first work that tackles the vision-brain problem with a continual learning approach (detailed in Fig. 1). Second, a new Subject Commonality (SC) module will be presented to capture shared brain activation patterns across subjects. Unlike prior methods, COBRA can automatically identify a broader range of common patterns. Third, we present a new Prompt-based Subject-Specific (PSS) module to capture unique patterns specific to each subject. The module is trained to form the most relevant prompt, representing each subject’s distinct characteristics. Fourth, we introduce a novel MRIFormer module that contains a transformer encoder-decode to construct fMRI features for VBU. Apart from prior methods, this module can explore the sequence information, flexible and independent of any length of target features, i.e., Contrastive Language-Image Pre-Training (CLIP) features. Fifth, we introduce a simple yet effective training strategy to address the CVBU problem. In particular, COBRA is only required to update PSS and MRIFormer modules for new subjects. It leaves previous subjects’ PSS and MRIFormer unaffected, thus addressing catastrophic forgetting. Notably, the PSS module training is label-free, eliminating the need for additional annotations. The extensive experiments demonstrate our method achieves state-of-the-art (SOTA) performance in VBU and continual learning. The pros and cons of COBRA compared to prior methods are summarized in Table 1.
@@ 2.1 Vision-Brain Understanding
In VBU, the fMRI signals corresponding to subjects’ viewing of an image are treated as input, which is then decoded to reconstruct the image originally viewed [27, 62, 63, 55, 72, 68, 77]. Many works have leveraged diffusion [67, 26, 11, 60, 78] to learn the fMRI signal distribution. MindEye [62] aligns the original image and fMRI latent spaces via contrastive learning and then uses diffusion to reconstruct the original image stimulus. MindEye2 [63] trains in a limited data setting to address expensive fMRI data acquisition. Takagi et al. [68] use latent diffusion [60] for image reconstruction, using a part of the fMRI as the diffusion conditioning. MindBridge [72] trains with a cyclic fMRI reconstruction mechanism, aligning brain data between subjects, allowing for more robust brain-to-image and text decoding using dual embeddings. Psychometry [55], by contrast, trains a singular model common and unique features for all subjects. NeuroPictor [27] uses a latent diffusion-like approach [60] to encode the fMRI signals into a latent space to separate high- and low-level features, giving the ability to focus on finer details. UMBRAE [77] decodes fMRI signals in a vision-language framework, aligning text descriptions and image features.
@@ 2.2 Continual Learning
We review two continual learning approaches [73], each with advantages and limitations. Rehearsal-based methods address catastrophic forgetting by storing sampled past tasks’ data in buffers [25, 6, 7, 5, 2, 58, 66, 54, 4, 76], and thus integrate old and new data to prevent knowledge loss. Several works leverage knowledge distillation to compress knowledge across older and newer tasks [7, 2, 58, 76] while others leverage self-supervised methods [54, 4, 42, 52, 43, 44, 41, 45, 48, 47, 46, 49, 39, 51, 40, 50]. Key limitations include buffer size, where making too small degrade learning [4] and data privacy, restricting access to past data [66]. Architecture-based methods address catastrophic forgetting by modifying the architecture itself. Typically, this is done by adding a new set of parameters for new tasks [81, 79, 61, 57, 35, 29] or sub-networks specialized for a certain task are maintained [75, 64, 37, 28]. Key limitations include the added complexity of added parameters and the model needing the task type beforehand to select the task parameters, which is not guaranteed to be given at test time.
@@ 2.3 Limitations of Data Representation in Prior Methods
L1. Since the brain sizes and structures of each subject differ, the sizes of fMRI signals vary across subjects. These signal sizes range from approximately 13,000 to 16,000 voxels per subject. Prior methods [10, 80, 31, 69, 3, 24, 68, 62, 55, 72, 63, 27] fixed the input signal size and adapted to new subjects with different input sizes. This poses a significant challenge to continual brain understanding, requiring methods to work with inputs with varying dimensions.
@@
L3. Prior methods rely on signals from specific Regions of Interest (ROI), such as floc-face, floc-body, and floc-place, rich in information about human visual perception. However, this focus may overlook other neighborhoods’ information. For instance, the floc-face area activates when a subject views a face in a visual stimulus, while the floc-body is responsible for human body perception. This approach might fail while facing with non-human subjects. To the best of our knowledge, no specific ROIs are dedicated to such common objects. Therefore, using signals from only the face, body, place, and word ROIs might miss critical information related to other elements in the visual stimulus.
@@ 3.1 Continual Vision-Brain Understanding
We define a sequence of brain understanding tasks: , where the subject is a set of pairs where each pair contains a visual stimulus and its corresponding record fMRI signals . is the number of samples for the task and is the sample index of task dataset. The goal is to train a model to predict for any unseen sample from arbitrary subjects. In addition, the data of previous subjects are not available in the training process of new subjects.
@@ 3.2 Data Preprocessing
To address the limitations of L1, we adopt the 3D FreeSurfer [18] (fsaverage) space to represent the fMRI signals to the 2D forms. Since the 3D FreeSurfer is a general template of the human brain, it can effectively represent the fMRI signals of all subjects. In addition, the 3D FreeSurfer template offers 3D spatial information, which can help to address L2. While L3 will be majorly addressed via our Subject Commonality Module (presented in the next section), we have found that the data preprocessing step plays an important role in providing the essential conditions to address L3. In particular, instead of using fMRI signals from specific ROIs, we use the entire brain signal, hypothesizing that the model will learn to identify the relevant regions for each ROI, automatically. In conclusion, representing brain signals in 2D addresses the previous limitations and offers flexibility for various tasks like continual brain understanding and unified model learning. It provides an advanced approach to understanding the brain by leveraging CNNs or transformers. Fig. 2 (right-hand side) illustrates the representation of the 2D form of the fMRI signals. For convenience, in the rest of the paper, we refer to fMRI signals in the 2D flattened form instead of in 1D.
@@ 3.3 Subject Commonality Module
This section introduces the Subject Commonality (SC) module, designed to capture shared features across subjects. This module is inspired by neuroscience research to identify the functional localizers (floc) of brain patterns related to specific visual stimuli [13, 9, 15, 74, 16, 17, 14, 23, 22, 20]. This module also can address L3 and discover ROIs of different categories rather than face, body, etc. Given the fMRI signals represented in 2D format, we define SC module as a transformer-based network, denoted as , to extract the commonality feature where contains high-level patch tokens which are output of Patch Embedding module, is number of patches. denoted as CLS token. The goal of this module is to find out what objects fMRI signals are representing. The output of the SC module is formed as follows,
where is a linear layer, is the predicted objects in the fMRI signals, is number of object. In the Natural Scenes Dataset [1] (NSD), is fixed to 80 since all stimuli are from the COCO [33] database. In addition, each stimulus contains multiple objects, for that reason, the is optimized using Binary Cross Entropy as in the Eqn (2).
Since is designed to receive the signal from any subjects, will learn common representation across subjects. Secondly, aims to learn the context inside the signals, i.e., what objects that signal contains.
@@ 3.4 Prompt-based Subject Specific (PSS) Module
To achieve this goal, we draw inspiration from recent advances in prompt-based learning or prompting [34], a new transfer learning technique in natural language processing (NLP). Prompting techniques design model textual inputs with templated or learnable prompt tokens containing additional task-specific information. In this paper, we treat prompts as subject-specific features. We need to form a prompt that describes that unique.
Methods. First, we need to build a pool of prompt tokens where these tokens contain semantic information of fMRI signals. Fortunately, we found that the high-level tokens can fulfill these requirements. Let be prompt tokens where is a linear layer. Let be a set of keys where each key will be associated with a token. To construct a prompt, we design the key-query paradigm to select the most -suitable tokens w.r.t as in the Eqn (3).
where denotes a matrix multiplication operation, is the subject-specific feature. We feed into a pooling followed by a linear layer and then classify the feature corresponding to the right subjects as in Eqn (4).
@@ 3.5 MRIFormer
The feature will be aligned with the corresponding Contrastive Language-Image Pre-Training (CLIP) feature of the visual stimulus: . We observe an inconsistency in the length of the feature, i.e., . The prior studies [68, 62, 72, 63] used a linear layer to map the fMRI feature to the same length as the CLIP one. However, it will lose the sequence properties of a feature. To address this problem, we take inspiration from Neural Machine Translation to design a transformer decoder to translate fMRI features to CLIP features. This approach can preserve sequential information and address feature length inconsistency as in Eqn (7).
@@ 3.7 Implementation Details
The 2D fMRI signals are resized to . We select vit-base-16 as the SC module to extract patches’ feature of the fMRI signals. Each feature vector has dimension . In the PSS module, we design associated keys for prompt tokens as a vector of . We select as a top-k number. That means for each query of commonality feature , we select top-30 tokens from the high-level tokens to describe specific features of the subject. The transformer encoder in Eqn (6) and the decoder in Eqn. (7) are designed to have four in-depth, dimension dim is 768, and the number of the head is set to 12. The framework is implemented via Pytorch and trained by 16 A100 GPUs (40GB each). The learning rate is set to initially and then reduced to zero gradually under ConsineLinear [38] policy. The batch size is set to /GPU. The weight factors are set to one equally in the Eqn 10. The model is optimized by AdamW [36] for epochs. The training is completed within two hours per subject, approximately.

### v3

@@ Abstract
Vision-Brain Understanding (VBU) aims to extract visual information perceived by humans from brain activity recorded through functional Magnetic Resonance Imaging (fMRI). Despite notable advancements in recent years, existing studies in VBU continue to face the challenge of catastrophic forgetting, where models lose knowledge from prior subjects as they adapt to new ones. Addressing continual learning in this field is, therefore, essential. This paper introduces a novel framework called Continual Learning for Vision-Brain (COBRA) to address continual learning in VBU. Our approach includes three novel modules: a Subject Commonality (SC) module, a Prompt-based Subject Specific (PSS) module, and a transformer-based module for fMRI, denoted as MRIFormer module. The SC module captures shared vision-brain patterns across subjects, preserving this knowledge as the model encounters new subjects, thereby reducing the impact of catastrophic forgetting. On the other hand, the PSS module learns unique vision-brain patterns specific to each subject. Finally, the MRIFormer module comprises a transformer encoder and decoder that learn the fMRI features for VBU from both common and specific patterns. In a continual learning setup, COBRA is trained on new PSS and MRIFormer modules for new subjects, while the modules for previous subjects remain unaffected. As a result, COBRA effectively addresses catastrophic forgetting and achieves state-of-the-art performance in both continual learning and vision-brain reconstruction tasks, surpassing previous methods.
keywords:
Continual Learning, Vision Brain Understanding, Neurocomputing, fMRI, Reconstruction1 Introduction
@@
Recently, vision-brain understanding (VBU) has made substantial progress due to advances in generative models, including GANs [2, 3] and Diffusion models [4, 5, 1, 6, 7, 8]. These approaches allow the reconstruction of increasingly realistic and accurate images from functional Magnetic Resonance Imaging (fMRI) signals. The approaches designed per-model-per-subject, as seen in [4, 5], are required to train a new model for a new subject. It can learn subject-specific but ignore subject-commonality features. To address this limitation, the studies [1, 6, 7, 8] proposed unified models trained on multiple subjects simultaneously. While these approaches are trained on a close set of subjects, they may still face a degradation in performance on new subjects. Recent efforts [1, 8] have aimed to adapt pre-trained models to new subject data, yet they continue to suffer from performance drops on prior subjects due to catastrophic forgetting [9, 10, 11], where models lose learned knowledge when adapting to new subject shifts. The ideal solution to this problem is to update the entire model with prior and new data. However, given the sensitive nature of brain data and the importance of privacy, retaining previous training data is often infeasible. Moreover, as the number of subjects grows, computational training costs become a concern. Thus, in practice, VBU models need to continually learn from new subjects without requiring retraining on prior data. This paradigm is known as Continual Vision-Brain Understanding (CVBU). Given the limitation in previous brain understanding approaches, we raise critical questions for CVBU: (1) What types of knowledge can the model retain, and (2) what types of knowledge must it learn specifically from new subject data?
@@
Since all subjects share similar patterns of brain activations [12, 13, 14, 15, 16, 17, 18, 19, 20, 21], if the model can understand these patterns, there is no need to learn from the data of new subjects. Consequently, the answer to the first question is to retain the knowledge of common patterns of brain activations learned from prior subjects and apply them to new ones. To achieve this, we propose a novel module designed to capture these patterns. Unlike prior methods [12, 13, 14, 15, 16, 22, 23, 24, 17, 18, 19, 20, 21], our proposed approach can automatically identify a broader set of common patterns.
@@
Contributions of this Work: In summary, this work presents a novel continual learning approach to VBU. The contributions of this work can be summarized as follows. First, we introduce a new Continual Learning for Vision-Brain Understanding (COBRA) framework to the visual-brain understanding problem. To the best of our knowledge, this is the first work to tackle the vision-brain problem using a continual learning approach (detailed in Fig. 1). Second, a new Subject Commonality (SC) module will be presented to capture shared brain activation patterns across subjects. Unlike prior methods, COBRA can automatically identify a broader range of common patterns. Third, we present a new Prompt-based Subject-Specific (PSS) module to capture unique patterns specific to each subject. The module is trained to form the most relevant prompt, representing each subject’s distinct characteristics. Fourth, we introduce a novel MRIFormer module that contains a transformer encoder-decoder to construct fMRI features for VBU. Apart from prior methods, this module can explore sequence information, being flexible and independent of the length of target features, i.e., Contrastive Language-Image Pre-Training (CLIP) features. Fifth, we introduce a simple yet effective training strategy to address the CVBU problem. In particular, COBRA is only required to update the PSS and MRIFormer modules for new subjects. It leaves previous subjects’ PSS and MRIFormer unaffected, thus addressing catastrophic forgetting. Notably, the PSS module training is label-free, eliminating the need for additional annotations. The extensive experiments demonstrate that our method achieves state-of-the-art (SOTA) performance in VBU and continual learning. The pros and cons of COBRA compared to prior methods are summarized in Table 1.
@@ 2.1 Vision-Brain Understanding
In VBU, the fMRI signals corresponding to subjects’ viewing of an image are treated as input, which is then decoded to reconstruct the image originally viewed [7, 5, 8, 6, 1, 4, 26]. Many works have leveraged diffusion [27, 28, 29, 30, 31] to learn the distribution of the fMRI signal. MindEye [5] aligns the original image and fMRI latent spaces via contrastive learning and then uses diffusion to reconstruct the original image stimulus. MindEye2 [8] trains in a limited data setting to address expensive fMRI data acquisition. Takagi et al. [4] use latent diffusion [30] for image reconstruction, using a part of the fMRI as the diffusion conditioning. MindBridge [1] trains with a cyclic fMRI reconstruction mechanism, aligning brain data across subjects, which enables more robust brain-to-image and text decoding using dual embeddings. Psychometry [6], by contrast, trains a singular model with common and unique features for all subjects. NeuroPictor [7] uses a latent diffusion-like approach [30] to encode the fMRI signals into a latent space to separate high- and low-level features, giving the ability to focus on finer details. UMBRAE [26] decodes fMRI signals in a vision-language framework, aligning text descriptions and image features.
@@ 2.2 Continual Learning
We review two continual learning approaches [32], each with advantages and limitations. Rehearsal-based methods address catastrophic forgetting by storing sampled past tasks’ data in buffers [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], and thus integrate old and new data to prevent knowledge loss. Several works leverage knowledge distillation to compress knowledge across older and newer tasks [35, 37, 38, 42] while others leverage self-supervised methods [40, 41, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Key limitations include buffer size, where making it too small degrades learning [41] and data privacy, restricting access to past data [39]. Architecture-based methods address catastrophic forgetting by modifying the architecture itself. Typically, this is done by adding a new set of parameters for new tasks [57, 58, 59, 60, 61, 62] or sub-networks specialized for a certain task are maintained [63, 64, 65, 66]. Key limitations include the added complexity of added parameters and the model needing the task type beforehand to select the task parameters, which is not guaranteed to be given at test time.
@@ 2.3 Limitations of Data Representation in Prior Methods
L1. Since the brain sizes and structures of each subject differ, the sizes of fMRI signals vary across subjects. These signal sizes range from approximately 13,000 to 16,000 voxels per subject. Prior methods [67, 68, 69, 70, 71, 72, 4, 5, 6, 1, 8, 7] fixed the input signal size and adapted to new subjects with different input sizes. This poses a significant challenge to understanding the brain continually, requiring methods to work with inputs of varying dimensions.
@@
L3. Prior methods rely on signals from specific Regions of Interest (ROI), such as floc-face, floc-body, and floc-place, rich in information about human visual perception. However, this focus may overlook information about other neighborhoods. For instance, the floc-face area activates when a subject views a face in a visual stimulus, while the floc-body is responsible for perceiving the human body. This approach may fail when applied to non-human subjects. To the best of our knowledge, no specific ROIs are dedicated to such common objects. Therefore, using signals from only the face, body, place, and word ROIs might miss critical information related to other elements in the visual stimulus.
@@ 3.1 Continual Vision-Brain Understanding
We define a sequence of brain understanding tasks: , where the subject is a set of pairs where each pair contains a visual stimulus and its corresponding record fMRI signals . is the number of samples for the task, and is the sample index of the task dataset. The goal of subject incremental learning in VBU is to train a model to predict for minimizing the total risk of all subjects.
where calculates the expected loss error of while training with subject . is the total error evaluated on the past subjects trained on the previous steps. The challenge in the CVBU is that data points from previous subjects may not be available during training on new subjects due to several reasons, such as privacy concerns. This paper aims to tackle this problem by learning from new subjects while maintaining the performance of past subjects. In traditional task-incremental continual learning, tasks differ by objective (e.g., classification of different categories), however, in CVBU, each subject shares the same objective, e.g., reconstructing or decoding the same image space. The CVBU setup, however, introduces a unique distribution shift due to individual brain structure, neural responses, and voxel patterns. This makes each subject a non-independent and identically distributed (non-i.i.d. distribution) over the same task space rather than a new task. Moreover, CVBU has unique constraints: (1) no access to past subject data due to privacy, (2) high-dimensional, structured, and spatially variable input (fMRI), and (3) shared decoder space (CLIP features), unlike task-specific heads used in task-incremental CL. The details of the proposed method will be presented in the next sections.
@@ 3.2 Data Preprocessing
To address the limitations of L1, we adopt the 3D FreeSurfer [73] (fsaverage) space to represent the fMRI signals in the 2D forms. Since the 3D FreeSurfer is a general template of the human brain, it can effectively represent the fMRI signals of all subjects. In addition, the 3D FreeSurfer template offers 3D spatial information, which can help to address L2. While L3 will be majorly addressed via our Subject Commonality Module (presented in the next section), we have found that the data preprocessing step plays an important role in providing the essential conditions to address L3. In particular, instead of using fMRI signals from specific ROIs, we use the entire brain signal, hypothesizing that the model will learn to identify the relevant regions for each ROI, automatically. In conclusion, representing brain signals in 2D addresses the previous limitations and offers flexibility for various tasks, such as continuous brain understanding and unified model learning. It provides an advanced approach to understanding the brain by leveraging CNNs or transformers. Fig. 2 (right-hand side) illustrates the representation of the 2D form of the fMRI signals. It is important to note that, the 2D form of fMRI signals does not lead to information loss since data transformations are just purely projection operations. In particular, the 2D representation used in COBRA is derived from flattening the 3D cortical surface (fsaverage) into 2D, preserving geodesic distance and spatial relationships, similar to how FreeSurfer is used in neuroscience. Importantly, this is a bijective projection. Each voxel has a unique mapping from 3D to 2D and vice versa, with no downsampling or interpolation. Thus, the original spatial resolution is maintained. To further support this, in the Appendix Sec 7.2, we explain more about this issue. For convenience, in the rest of the paper, we refer to fMRI signals in the 2D flattened form instead of in 1D.
@@ 3.3 Subject Commonality Module
This section introduces the Subject Commonality (SC) module, designed to capture shared features across subjects. This module is inspired by neuroscience research to identify the functional localizers (floc) of brain patterns related to specific visual stimuli [12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. This module can also address L3 and discover ROIs of different categories rather than face, body, etc.
Given the fMRI signals represented in 2D format, we define SC module as a transformer-based network, denoted as , to extract the commonality feature where contains high-level patch tokens which are output of Patch Embedding module, is number of patches. denoted as CLS token. The goal of this module is to find out what objects fMRI signals represent. The output of the SC module is formed as follows,
where is a linear layer, is the predicted objects in the fMRI signals, is number of object. In the Natural Scenes Dataset [74] (NSD), is fixed to 80 since all stimuli are from the COCO [75] database. In addition, each stimulus contains multiple objects; for that reason, the is optimized using Binary Cross Entropy as in the Eqn (3).
Since is designed to receive the signal from any subjects, will learn common representation across subjects. Secondly, aims to learn the context inside the signals, i.e., what objects the signal contains.
@@ 3.4 Prompt-based Subject Specific (PSS) Module
To achieve this goal, we draw inspiration from recent advances in prompt-based learning, also known as prompting [25, 76, 77], a novel transfer learning technique in natural language processing (NLP). Prompting techniques design a model to generate textual inputs with templated or learnable prompt tokens containing additional task-specific information. In this paper, we treat prompts as subject-specific features. We need to form a prompt that describes this unique aspect.
Methods. First, we need to build a pool of prompt tokens that contain the semantic information of fMRI signals. Fortunately, we found that the high-level tokens can fulfill these requirements. Let be prompt tokens where is a linear layer. Let be a set of keys where each key will be associated with a token. To construct a prompt, we design the key-query paradigm to select the most -suitable tokens w.r.t as in the Eqn (4).
where denotes a matrix multiplication operation, is the subject-specific feature. We feed into a pooling layer followed by a linear layer and then classify the feature corresponding to the right subjects as in Eqn (5).
@@ 3.5 MRIFormer
The feature will be aligned with the corresponding Contrastive Language-Image Pre-Training (CLIP) feature of the visual stimulus: . We observe an inconsistency in the length of the feature, i.e., . Prior studies [4, 5, 1, 8] utilized a linear layer to map the fMRI feature to the same length as the CLIP one. However, it will lose the sequence properties of a feature. To address this problem, we take inspiration from Neural Machine Translation to design a transformer decoder to translate fMRI features to CLIP features. This approach can preserve sequential information and address feature length inconsistency, as shown in Eqn. (8).
@@ 3.7 Implementation Details
The 2D fMRI signals are resized to . We select vit-base-16 as the SC module to extract the patches’ features of the fMRI signals. Each feature vector has dimension . In the PSS module, we design associated keys for prompt tokens as a vector of . We select as the top number. That means for each query of commonality feature , we select top-30 tokens from the high-level tokens to describe specific features of the subject. The transformer encoder in Eqn (7) and the decoder in Eqn. (8) is designed to have four in-depth, dimension dim is 768, and the number of heads is set to 12. The framework is implemented using PyTorch and trained on 16 A100 GPUs (40GB each). The learning rate is set to initially and then reduced to zero gradually under ConsineLinear [79] policy. The batch size is set to /GPU. The weight factors are set to one equally in the Eqn 11. The model is optimized by AdamW [80] for epochs. The training is completed within approximately two hours per subject.